# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

### v1

@@ 4.1 Iterative formulation of JMA
  • When the target class is reached, a binary search is performed between the perturbed sample $x+\delta^{*}$ and the original input $x$, and the adversarial example resulting in the smallest perturbation is taken as the final output of the attack.
Finally, in practice, after that the optimal perturbation $\delta^{*}$ is superimposed to $x$, a clipping operation is performed to be sure that the solution remains in the $[0,1]^{m}$ range. Regarding the full rank of the Jacobian, this assumption has been verified experimentally and discussed in Section 5.4.
The iterative JMA attack resulting from the application of the above steps is described in Algorithm 1. In the algorithm, the maximum number of iterations (updates) is set to $n_{it,max}$.
Algorithm 1 Iterative Jacobian-induced MAhalanobis distance (JMA) attack
@@
4:     calculate matrix $A$ and vector $b$
@@
4.2 Simplified formulation for the multi-label case
@@ 5 Experimental Setting
To evaluate the performance of JMA, we have run several experiments addressing different classification scenarios, including ECOC-based classification [9] and multi-label classification. Although less significant for our study, we have also run some experiments in the one-hot encoding scenario. In all the cases, the JMA attack is applied at the logits level.
@@
We implemented the JMA attack by using Python 3.6.9 via the Keras 2.3.1 API. We run the experiments by using an NVIDIA GeForce RTX 2080 Ti GPU. The code, as well as the trained models and all the information for the reproducibility of the experiments are publicly available at https://github.com/guowei-cn/JMA–A-General-Close-to-Optimal-Targeted-Adversarial-Attack-with-Improved-Efficiency.git.
In the following sections, we describe the classification tasks considered in the experiments, present the evaluation methodology and comparison analysis, and detail the experimental setting.
5.1 Classification networks and settings
5.1.1 One-hot encoding
To assess the performance of JMA attack in the one-hot encoding scenario, we considered traffic sign classification on the German Traffic Sign Recognition Benchmark (GTSRB) dataset [25], with a VGG16 architecture. The error rate of the trained models on clean images is 0.005.
5.1.2 ECOC-based classifier
@@
We considered the ECOC framework for three image classification tasks, namely MNIST classification [26], CIFAR-10 classification [27], and GTSRB classification [25]. We implemented the ECOC scheme considering an ensemble of networks, each one outputting one bit [^6] According to [9], resorting to an ensemble of networks permits to achieve better robustness against attacks.. The number of branches $h$ of the ensemble is $h=10$ for MNIST and CIFAR-10, and $h=16$ for GTSRB. We used Hadamard codewords as suggested in [9]. We considered an Hadamard code with $n=16$ for MNIST and CIFAR-10. For the GTSRB case, we set $l=32$, by selecting the classes with more examples, and used an Hadamart code with $n=32$. Following the original ECOC design, we considered the VGG16 architecture [28] as baseline, with the first 6 convolutional layers shared by all the networks of the ensemble (shared bottom), and the remaining 10 layers (the last 8 convolutional layers and the 2 fully connected layers) that have been trained separately for each ensemble branch. The weights of the Image-Net pre-trained model are used for the shared bottom.
5.1.3 Multi-label classification
For the multi-label experiments, we considered the model architecture adopted in [10] that is, a standard InceptionV3 [29]. The categorical hinge loss was used to train the model, with a standard Adam optimizer with lr=$10^{-4}$ and batch size 64.
TABLE I: Types of objects in the VOC2012 dataset.
Category Object name
Person person
Animal bird, cat, cow, dog, horse, sheep
Vehicle aeroplane, bicycle, boat, bus, car, motorbike, train
Indoor bottle, chair, dining table, potted plant, sofa, tv/monitor
@@ 5.1.3 Multi-label classification
The model was trained on the VOC2012 dataset [30], which is a benchmark dataset adopted in several multi-class classification works [14, 31]. This dataset has been used to train models to recognize object inside scenes. Specifically, the 11,530 images of the dataset contain objects from 20 possible object classes, see Table I (the objects are classified into four categories). Then, $n=20$ and the number of possible label vectors is $2^{n}\approx 10^{6}$. All datasets are split into training, validation, and testing subsets, with proportion 6:2:2. The performance of the network in this case are measured in terms of mean average precision (mAP), that is, the mean of the average precision (AP) for all the classes, where AP is a measure of the area under the precision-recall curve. According to our experiments, the mAP of our model is 0.93, which is aligned with [10].
5.2 Comparison with the state-of-the-art
@@
We have compared the performance of JMA with those of the most relevant state-of-the-art attacks for each classification scenario. In all the cases, we used the codes made available by the authors in their repositories.
@@
For the multi-label case, we considered ML-C&W [10], MLA-LP [11] and LOTS [18] (we have not considered ML-DF as its performance are always inferior to ML-C&W, see [10, 11]). As described in Section 3, LOTS can be applied to any internal layer of the network. In our experiments, for a fair comparison, we have applied LOTS at the logits layer. By following [18], for every target label vector, we randomly selected 20 images with that label vector, and averaged the logits to get the target output vector (for target label vectors for which the total number of images available is lower than 20, we used all the images). We stress that, in the way it works, LOTS can not target an arbitrary labels’ combination, since examples for that combination may not be available.
@@
For the case of ECOC classification, the comparison is carried out against C&W, which is also the attack considered in [9], and [17] (ECOC Attack) specialized to work with the ECOC model [17]. Although suboptimal in this case, we also considered ML-C&W and MLA-LP, by applying the latter as discussed in Section 3.2. However, given the complexity of the attack in the ECOC scenario and the high sub-optimality of MLA-LP, the performance of this algorithm are extremely poor, and no image can be attacked by using this method. Finally, the performance of LOTS attack are also assessed. In this case, the target output vectors are computed by averaging on 50 images (when the number of available images is lower than 50, we averaged all of them).
@@
5.3 Evaluation methodology
@@
To test the attacks in the various scenarios, we proceeded as follows: we randomly picked 200 images from the test set (among those that are correctly classified by the network model); then, for each image, we randomly picked a target label vector, different from the true label vector.
@@
For multi-label classification, in principle, there can be - and in fact there are - combinations of labels that are not represented in the training set. it is also possible that some of these combinations do not correspond to valid label vectors, as some labels may not be simultaneously 1 (or 0). Since LOTS requires the availability of samples belonging to the target label vector, for ease of comparison, in our experiments, we considered label vectors that appear in the training set as target label vectors. For the other methods, some tests have also been carried out in the more challenging case where the target label vectors correspond to randomly chosen target vectors. Such target vectors are obtained by randomly changing a prescribed number of bits in the label vector of the to-be-attacked image.
@@
An attack is considered successful only when the predicted label vector and the target label vector coincide. If only a subset of the labels can be modified by the attack, and the target label vector is not reached, we mark this as a failure. We measured the performance of the attacks in terms of Attack Success Rate (ASR), namely, the percentage of generated adversarial examples that are assigned to the target class. To measure the quality of the adversarial image, we considered the Peak Signal-Noise Ratio (PSNR), characterizing the distortion introduced by the attack. The PSNR is evaluated as $20\log_{10}(255\cdot\sqrt{H\times W}/||\delta||_{2})$, where $H\times W$ is the size of the image [^7] Given that the attacks are multi-step and iterative, and different iterations are necessary for the various algorithms, we can not use $\epsilon$ as an indication of the distortion.. For the tests with random choice of the target vector (for a prescribed number of flipped labels), we also report the average percentage of labels/bits successfully modified by the attack, indicated as bASR.
@@
The computational complexity is measured by providing the average number of iterations required by the attack, and the time - in seconds - necessary to run it. In the following, we let $n_{it}$ denote the number of iterations necessary to run an attack, or the number of attack updates. The average number of iterations is denoted by $\bar{n}_{it}$.
5.4 Setting the Attack Parameters
@@
The tunable parameters of JMA are the maximum number of iterations (${n}_{it,max}$) and the step size $\epsilon$. The iteration parameter takes also into account the number of ’for’ loops and the number of steps of the binary search in the final iteration (see Algorithm 1), the total number of updates of JMA, then, is ${n}_{it}=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops of the algorithm ($v<{n}_{it,max}$, see Algorithm 1) and $n_{bs}$ is the number of steps of the binary search. $n_{bs}$ is set to 6 in our the experiments.
For C&W, ML-C&W and ECOC attack, the parameters are the initial $\lambda$, the number of steps of the binary search $n_{bs}$, and the maximum number of iterations $n_{it,max}$. All these attacks carry out a binary search over the parameters $\lambda$ at every iteration to find the best value. Then, $n_{it,max}$ provides a constraint on $n_{bs}\cdot b$, where $b$ is the number of iterations of the optimization algorithm. As for their setting, we consider $n_{bs}=5$ and $10$, and various values of $\lambda$.
@@
Finally, for MLA-LP and LOTS, the only parameter is the maximum number of iterations $n_{it,max}$ allowed in the gradient descent. In our experiments, we considered various values of $n_{it,max}$. For MLA-LP, we set $n_{it,max}=100$ for the multi-label classification case, and raised it up to $3000$ in the ECOC case when the attack is more difficult. For LOTS, we set $n_{it,max}=2000$, namely the maximum number considered in [18], maximizing the chances to find an adversarial example. Given that the time complexity of LOTS is limited compared to the other comparison methods, in fact, the time complexity of the attack remains small also for large values of $n_{it,max}$. We have verified experimentally that LOTS converges within 1000 iterations 92% of the times [^8] Convergence is determined by checking whether the new loss value is close enough to the average loss value of the last 10 iterations., validating the soundness of our choice. For ML-C&W, we set $n_{it,max}=1000$ as default value, and in most of the cases results are reported for the setting $(\lambda,n_{bs},n_{it,max})=(0.01,10,1000)$, that resulted in the highest ASR. Finally, for JMA, we generally got no benefit by increasing the number of iterations beyond 200, since most of the time JMA converges in few iterations, and 200 is more than enough to get and adversarial image in all the cases.
@@
For sake of simplicity and comparison, the confidence parameter is set to 0 for all the attack methods. To avoid that the an attack gets stuck with some images (e.g. in the multi-label case when many labels have to be changed), we set a limit on the running time of the attack, and considered the attack unsuccessful when an adversarial example could not be generated in less than 1 hour.
@@ 6 Results
The results we obtained in the various classification settings are reported and discussed in this section. In all the tables, $\bar{n}_{it}$ refers to the average number of iterations/updating of the attack, averaged on the successful attacks. The attack parameters we report are $(\lambda,n_{bs},n_{it,max})$ for C&W, ML-C&W and ECOC attack, and $(\epsilon,n_{it,max})$ for JMA.
6.1 One-hot encoding
@@
As we said, the main advantage of JMA with respect to state of the art methods is obtained for classifiers that do not adopt one-hot encoding. In such a scenario, in fact, focusing only on the target logit or final score, and ignoring the others, as done by the most common adversarial attacks, is approximately optimal. Nevertheless, we verified that the JMA method is also effective in the one-hot encoding scenario. Table II reports the results we got on the GTSRB dataset. We see that JMA achieves a much higher ASR compared to C&W, getting close to 1, yet with a lower PSNR. This is likely due to the fact that JMA tends to modify all the output scores, while C&W directly focuses on raising the target output, which is an effective strategy when the one-hot scheme is adopted for the encoding. In addition, JMA can obtain an adversarial image in a much smaller number of iterations, requiring only very few seconds (2.45 sec in the best setting) to attack an image thus resulting much faster than C&W attack.
TABLE II: Results of the attacks against an one-hot encoding classifier for the GTSRB dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA ( 0.01, 200) 0.45 43.16 66.49 9.82
(0.05, 200) 0.81 37.88 50.65 7.44
(0.1, 200) 0.88 36.86 38.00 5.66
(0.2, 200) 0.94 35.76 29.63 4.41
(0.6,200) 0.98 33.67 16.20 2.46
C&W (1e-4,5, 100) 0.19 46.65 85.95 45.06
(1e-4,5,200) 0.33 45.24 171.88 90.11
(1e-2,5,500 ) 0.55 42.36 443.18 233.45
(1e-1,10,2000) 0.75 41.14 1661.11 875.35
(1e-1,10,5000) 0.82 41.02 3577.5 1884.59
6.2 ECOC-based classifiers

### v2

@@ V-A Iterative formulation of JMA
  • When the target class is reached, a binary search is performed between the perturbed sample $x+\delta^{*}$ and the original input $x$, and the adversarial example resulting in the smallest perturbation is considered.
Finally, in practice, after that the optimal perturbation $\delta^{*}$ is superimposed to $x$, a clipping operation is performed to be sure that the solution remains in the $0,1]^{m}$ range.
The iterative version of JMA resulting from the application of the above steps is described in Algorithm 1. In the algorithm, the maximum number of iterations (updates) is set to $n_{it,max}$.
Algorithm 1 Jacobian-induced Mahalanobis-distance attack
@@
4:  calculate matrix AA and vector $b$
@@
V-B Simplified formulation for the multi-label case
@@ VI Experimental Setting
To evaluate the performance of JMA, we have run several experiments addressing different classification scenarios, including ECOC-based classification [8] and multi-label classification. Although less significant for this study, we have also run some experiments in the one-hot encoding scenario. In all the cases, JMA is applied at the logits level.
@@
We implemented JMA by using Python 3.6.9 via the Keras 2.3.1 API. We run the experiments by using an NVIDIA GeForce RTX 2080 Ti GPU. The code, as well as the trained models and the information for the reproducibility of the experiments are publicly available at https://github.com/guowei-cn/JMA–A-General-Close-to-Optimal-Targeted-Adversarial-Attack-with-Improved-Efficiency.git. In the following sections, we describe the classification tasks considered in the experiments, present the evaluation and comparison methodologies, and detail the experimental setting.
VI-A Classification networks and settings
VI-A1 One-hot encoding
To assess the performance of JMA in the single-label classification scenario, we considered the task of traffic sign classification on the German Traffic Sign Recognition Benchmark (GTSRB) dataset [31], with a VGG16 architecture. In this scenario, the accuracy of the the trained model on clean images is 0.995. Moreover, to validate the effectiveness of the JMA algorithm in a challenging classification scenario with a large number of classes, we defined a new classification task by selecting 2,000 classes from the most populated categories in the ImageNet21K dataset [32] (referred to as ImageNet2K in the following). A ViT-B/16 network pretrained on ImageNet21K and fine-tuned on a training subset of ImageNet2K was used for this task. The accuracy on clean images is 0.66.
VI-A2 ECOC-based classifier
@@
We considered the ECOC framework for three image classification tasks, namely MNIST [33], CIFAR-10 [34], and GTSRB [31]. We implemented the ECOC scheme considering an ensemble of networks, each one outputting one bit [^6] According to [8], resorting to an ensemble of networks permits to achieve better robustness against attacks.. The number of branches $h$ of the ensemble is $h=10$ for MNIST and CIFAR-10, and $h=16$ for GTSRB. We used Hadamard codewords as suggested in [8]. We considered an Hadamard code with $n=16$ for MNIST and CIFAR-10. For the GTSRB case, we set $l=32$, by selecting the classes with more examples, and used an Hadamart code with $n=32$. Following the original ECOC design, we considered the VGG16 architecture [35] as baseline, with the first 6 convolutional layers shared by all the networks of the ensemble (shared bottom), and the remaining 10 layers (the last 8 convolutional layers and the 2 fully connected layers) trained separately for each ensemble branch. The weights of the ImageNet pre-trained model are used for the shared bottom.
VI-A3 Multi-label classification
@@
We primarily tested the performance of JMA in a multi-label setting by using the VOC2012 dataset [36], which is a benchmark dataset adopted in several multi-class classification works [24, 23]. This dataset has been used to train models to recognize object inside scenes. Specifically, the 11,530 images of the dataset contain objects from 20 classes. Then, $n=20$ and the number of possible label vectors is $2^{n}\approx 10^{6}$. The dataset is split into training, validation, and testing subsets, with proportion 6:2:2. We considered the same model architecture adopted in [9] that is, a standard InceptionV3 [37]. The categorical hinge loss was used to train the model, with a standard Adam optimizer with $lr=10^{-4}$ and batch size 64. The performance of the network in this case are measured in terms of mean average precision (mAP), that is, the mean of the average precision (AP) for all the classes, where AP is a measure of the area under the precision-recall curve. According to our experiments, the mAP of our model is 0.93, which is aligned with [9].
To better show the generality of the proposed attack, we also trained additional models using different and more modern architectures, namely, ResNet50 [38], ViT-B/16 [39], and CLIP+MLP [40] , and ran attacks against them The results regarding these networks are reported in Section VII-D. We further validated the effectiveness of JMA on multi-label classification by running some additional experiments on more complex datasets, namely MS-COCO [41] and NUS-WIDE [42] (see again Section VII-D).
VI-B Comparison with the state-of-the-art
@@
We have compared the performance of JMA with the most relevant state-of-the-art attacks for the various classification scenarios. In all the cases, we used the code made available by the authors in public repositories.
For the multi-label case, we considered ML-C&W [9], MLA-LP [10] , LOTS [21] and the method in [11], referred to as SemA-ML in the following. We did not consider ML-DF, since its performance are always inferior to those of ML-C&W [9, 10])
@@
Regarding LOTS, as described in Section IV, it can be applied to any internal layer of the network. In the experiments, for a fair comparison, we have applied LOTS at the logits layer. By following [21], for every target label vector, we randomly selected 20 images with that label vector, and averaged the logits to get the target output vector (for target label vectors for which the total number of images available is lower than 20, we used all of them). We stress that, in the way it works, LOTS can not target an arbitrary labels’ combination, since examples for that combination may not be available.
@@
For the case of ECOC-based classification, the comparison is carried out against C&W, which is also the attack considered in [8], and [20] (ECOC Attack) specialized to work with the ECOC model [20]. Although suboptimal in this case, we also considered ML-C&W and MLA-LP, by applying the latter as discussed in Section IV-B. However, given the complexity of the attack in the ECOC scenario and the high sub-optimality of MLA-LP, the performance of this attack are extremely poor, and no image can be attacked by using it. Finally, the performance of LOTS are also assessed. In this case, the target output vectors are computed by averaging 50 images (when the number of available images is lower than 50, we averaged all of them).
@@
We did not consider attacks like I-FGSM, PGD, and also AutoPGD and AutoAttack, as they cannot be applied to frameworks different than single-label classification (see discussion in Section II), e.g. to ECOC and in particular to the multi-label case, for which they would need a suitable extension via the definition of a proper loss function.
VI-C Evaluation methodology
@@
To test the attacks in the various scenarios, we proceeded as follows: we randomly picked 200 images from the test set, among those that are correctly classified by the network model (in the complex multi-label classification tasks of MS-COCO and NUS-WIDE, some label errors always occur, hence the attacked samples are chosen at random); then, for each image, we randomly picked a target label vector, different from the true label vector.
@@
For multi-label classification, in principle, there can be - and in fact there are - combinations of labels that are not represented in the training set. it is also possible that some of these combinations do not correspond to valid label vectors. Since LOTS requires the availability of samples belonging to the target label vector, for ease of comparison, in the experiments, we considered label vectors that appear in the training set as target label vectors. For the other methods, some tests have also been carried out in the more challenging case where the target label vectors correspond to randomly chosen target vectors. Such target vectors are obtained by randomly changing a prescribed number of bits in the label vector of the to-be-attacked image.
@@
An attack is considered successful only when the predicted label vector and the target label vector coincide. If only a subset of the labels can be modified by the attack, and the target label vector is not reached, we mark this as a failure. We measured the performance of the attacks in terms of Attack Success Rate (ASR), namely, the percentage of generated adversarial examples that are assigned to the target class. To measure the quality of the adversarial image, we considered the Mean Square Error (MSE) of the attack, evaluated as MSE $\delta||_{2}/\sqrt{H\times W\times 3}$, where $H\times W\times 3$ is the size of the image (image values are in the [0,1] range).
For the tests with random choice of the target vector, we also report the average percentage of labels/bits successfully modified by the attack, indicated as bASR.
@@
The computational complexity is measured by providing the average number of iterations required by the attack, and the time - in seconds - necessary to run it. In the following, we let $n_{it}$ denote the number of iterations necessary to run an attack, that is, the number of attack updates. The average number of iterations is denoted by $\bar{n}_{it}$.
VI-D Attack Parameters Setting
@@
The tunable parameters of JMA are the maximum number of iterations (${n}_{it,max}$) and the step size $\epsilon$. The maximum number of iterations takes also into account the number of ‘for’ loops and the number of steps of the binary search in the final iteration (see Algorithm 1), the total number of updates of JMA, then, is ${n}_{it}=(v-1)+n_{bs}$, where $v$ is the number of ‘for’ loops of the algorithm ($v<{n}_{it,max}$, see Algorithm 1) and $n_{bs}$ is the number of steps of the binary search. $n_{bs}$ is set to 6 in our the experiments.
@@
For C&W, ML-C&W and ECOC attack, the parameters are the initial $\lambda$, the number of steps of the binary search $n_{bs}$, and the maximum number of iterations $n_{it,max}$. For all these attacks, a binary search is carried out over the loss tradeoff parameter $\lambda$ (see Section IV-A). At each binary search step, iterations are run (until $n_{it,max}$) to obtain the adversarial example. The number of iterations that we report considers only the number of attack updates corresponding to the final $\lambda$ value used by the attack. Hence, in these cases, the actual complexity of the attack is better reflected by the attack time. As for the setting of these parameters, we consider $n_{bs}=5$ and $10$, and various values of the initial $\lambda$. For SemA-ML, we set $n_{it,max}=300$ following the official code, with the other settings left as the default. For MLA-LP and LOTS, the only parameter is the maximum number of iterations $n_{it,max}$ allowed in the gradient descent. In the experiments, we considered various values of $n_{it,max}$. For MLA-LP, we set $n_{it,max}=100$ for the multi-label classification case, and raised it up to $3000$ in the ECOC case when the attack is more difficult. For LOTS, we set $n_{it,max}=2000$, namely the maximum number considered in [21], maximizing the chances to find an adversarial example. Given that the time complexity of LOTS is limited compared to the other methods, in fact, the time complexity of the attack remains small also for large values of $n_{it,max}$. We have verified experimentally that LOTS converges within 1000 iterations 92% of the times [^7] Convergence is determined by checking whether the new loss value is close enough to the average loss value of the last 10 iterations., validating the soundness of our choice. For ML-C&W, we set $n_{it,max}=1000$ as default value, and in most of the cases results are reported for $(\lambda,n_{bs},n_{it,max})=(0.01,10,1000)$, that resulted in the highest ASR. Finally, for JMA, we generally got no benefit by increasing the number of iterations beyond 200, since most of the time JMA converges in by far less iterations, and 200 is more than enough to get and adversarial image in all the cases.
@@
To simplify the comparison, the confidence parameter is set to 0 for all the attacks. To avoid that the an attack gets stuck with some images (e.g. in the multi-label case when many labels have to be changed), we set a limit on the running time of the attack, and considered the attack unsuccessful when an adversarial example could not be generated in less than 1 hour.
@@ VII Results
The results we obtained in the various settings are reported in this section. In all the tables, $\bar{n}_{it}$ refers to the average number of iterations/updating of the attack, averaged on the successful attacks. The attack parameters reported are $(\lambda,n_{bs},n_{it,max})$ for C&W, ML-C&W and ECOC attack, $(\epsilon,n_{it,max})$ for JMA and $n_{it,max}$ for LOTS and MLA-LP.
VII-A One-hot encoding
The main advantage of JMA with respect to state of the art is obtained for classifiers that do not adopt one-hot encoding. In such a scenario, in fact, focusing only on the target logit or the final score, as done by the most common adversarial attacks, is approximately optimal. Nevertheless, we verified that JMA is also effective in the one-hot encoding scenario.
@@ VII-A One-hot encoding
Table I reports the results on GTSRB. We see that JMA achieves a larger ASR compared to C&W, yet with a larger MSE. This is due to the fact that JMA tends to modify all the output scores, while C&W directly focuses on raising the target output, which is an effective strategy in the one-hot encoding case. In addition, JMA can obtain an adversarial image with much lower iterations, requiring only very few seconds (2.45 sec in the best setting) to attack an image.
The results of JMA against ImageNet2K are reported in Table II. We observe that, although JMA is not specifically designed for this scenario, it can successfully attack all the images in very few iterations, with an MSE similar to that obtained on GTSRB. The attack time is larger, due to the higher complexity of the NNLS problem.
TABLE I: Results against one-hot encoding classification for GTSRB.
Parameters ASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.81 1.6e-4 50.65 7.44
(0.1, 200) 0.88 2.1e-4 38.00 5.66
(0.2, 200) 0.94 2.6e-4 29.63 4.41
(0.6,200) 0.98 4.3e-4 16.20 2.46
C&W (1e-4,5, 100) 0.19 2.2e-5 85.95 45.06
(1e-4,5,200) 0.33 3.0e-5 171.88 90.11
(1e-2,5, 500) 0.55 5.8e-5 443.18 233.45
(1e-1,10, 2000) 0.75 7.7e-5 1661.11 875.35
(1e-1,10, 5000) 0.82 7.9e-5 3577.5 1884.59
TABLE II: Results of JMA against one-hot encoding classification with a large number of classes (ImageNet2K).
Parameters ASR MSE $\bar{n}_{it}$ Time(s)
JMA (0.1, 200) 0.99 3.8e-6 26.59 445.61
(0.2, 200) 1.00 6.2e-6 19.17 297.00
(0.6, 200) 1.00 2.1e-5 13.59 178.78
(1, 200) 1.00 4.2e-5 12.11 160.06
VII-B ECOC-based classification